OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (-3,-7),m=-6/5

OpenStudy (anonymous):

y - y1 = m(x - x1) is your equation

OpenStudy (anonymous):

plug in your numbers and solve so that you get to y=mx+b form